### v3

# SURGE: On the Potential of Large Language Models as General-Purpose Surrogate Code Executors

Abstract
Neural surrogate models have emerged as powerful and efficient tools in data mining. Meanwhile, large language models (LLMs) have demonstrated remarkable capabilities in code-related tasks. We investigate a novel application: using LLMs as surrogate models for code execution prediction. Given LLMs’ unique ability to understand and process diverse programs, they present a promising direction for building general-purpose surrogate models. To systematically investigate this capability, we introduce SURGE, a comprehensive benchmark with problems covering key aspects: multi-language programming tasks, competition-level programming problems, repository-level code analysis, high-cost scientific computing, time-complexity-intensive algorithms, buggy code analysis, programs dependent on specific compilers or execution environments, and formal mathematical proof verification. Through extensive empirical analysis of open-source and proprietary LLMs, we examine scaling laws, data efficiency, and predictive accuracy. Our findings reveal important insights about the feasibility of LLMs as efficient surrogates for computational processes, with implications for automated software testing, program analysis, and computational resource optimization in data mining applications. Code and dataset are released at https://github.com/Imbernoulli/SURGE.
1 Introduction
Neural surrogate models (Zhang et al., 2024; Sun and Wang, 2019) have emerged as powerful tools in data mining and machine learning, efficiently approximating complex computational processes. Meanwhile, Large language models (LLMs) (Reid et al., 2024; Meta, 2024; Anthropic, 2024b; Hui et al., 2024; Bi et al., 2024) have demonstrated remarkable capabilities in code-related tasks (Lu et al., 2021a; Zheng et al., 2023; Luo et al., 2023; Team, 2024a; Guo et al., 2024), including code understanding (Ahmad et al., 2020; Chakraborty et al., 2020) and code generation (Li et al., 2018a; Parvez et al., 2018). However, an equally important yet underexplored question is whether LLMs can serve as general-purpose surrogate code executors, which predict the output and behavior of a program without actually running it. Recent study (Lyu et al., 2024) acknowledges its importance, however, it focuses on a case study rather than a systematic analysis.
The ability to predict code execution outcomes without execution has tremendous siginficance. In scientific computing, running simulations often requires substantial computational resources and is time-consuming, making it impractical to test every possible configuration (Lu and Ricciuto, 2019; Hesthaven and Ubbiali, 2018; Benner et al., 2015). In security-sensitive environments, executing untrusted code poses inherent risks, necessitating alternative mechanisms for assessing program behavior without exposing the system to potential vulnerabilities (Nebbione and Calzarossa, 2023; Shirazi et al., 2017; Wang et al., 2024). Additionally, some code requires highly specific execution environments, which may not always be available, making surrogate execution a valuable alternative (Queiroz et al., 2023; Gu et al., 2025). Moreover, accurately predicting a model’s potential outputs or errors is crucial for improving traditional tasks such as code understanding, code generation, and even math reasoning (Li et al., 2025). Lastly, many works use LLMs as reward models (RMs) in reinforcement learning. For code tasks, accurate execution prediction is key to a reliable RM (Ouyang et al., 2022).
Traditional approaches to surrogate code executing (King, 1976; Cadar and Sen, 2013) struggle to generalize across languages and suffer from scalability issues when applied to complex real-world codebases. Containerized environments (Merkel, 2014) mitigate dependency issues but still require full code execution. Recent efforts to train neural executors (Yan et al., 2020) focus on narrow tasks and lack the generality needed for real-world code. In contrast, LLMs’ capacity to internalize patterns from vast code corpora (Lu et al., 2021b; Chaudhary, 2023) suggests a path toward general-purpose surrogate code execution, though their limits remain unquantified.
Therefore, it’s crucial to understand the potential of LLMs as general-purpose surrogate code executors. In this paper, we try to figure out how well current LLMs function as general-purpose surrogate code executors and whether this ability can be improved through training.
To systematically investigate LLMs as GEneral-purpose SURrogate code executors, we introduce SURGE. It includes components: (1) fundamental programming tasks in multiple languages, (2) competition programming problems requiring deep logical inference, (3) repository-level codebases that test long-range dependencies, (4) scientific simulations and optimizations where direct execution is high-cost, (5) time-consuming logical algorithms that have high time-complexity, (6) buggy code that examines LLMs’ ability to predict runtime errors, (7) programs whose behavior depends on specific compiler versions or execution environments and (8) math theorem proving in formal language (LEAN4) (De Moura et al., 2015; Moura and Ullrich, 2021) which expects compilers to testify the proofs.
Through extensive empirical evaluation of 21 open-source and proprietary LLMs on SURGE, we provide the first large-scale study of LLMs’ capabilities as computational surrogates. Additionally, we investigate the impact of various factors including prompt engineering strategies, programming language characteristics, computational complexity, and execution time requirements on surrogate performance. Our findings reveal both the promising potential and current limitations of LLMs as general-purpose code execution surrogates. The performance of typical models on SURGE is shown in Figure 1.
Beyond evaluating existing models, we conduct a scaling law study to determine whether models’ performance improves with increased model size and data availability. Specifically, we train models with sizes ranging from {0.5/1.5/3/7}B on different scales of training data from the formal language subset of SURGE and analyze how model capacity and training data scale affect surrogate execution accuracy. Our scaling experiments demonstrate that performance consistently improves with both model size (from 0.5B to 7B parameters) and training steps, with larger models showing stronger learning capacity and higher performance ceilings throughout the training process (Figure 2).
In short, our work makes the following key contributions:
We introduce SURGE, the first but holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. It consists of subsets and problems.
We conduct an extensive evaluation of open-source and proprietary LLMs on SURGE, providing the first large-scale study assessing their capabilities.
We present a scaling law study with models of varying sizes and data of different scales, providing empirical insights into whether LLMs systematically improve with more capacity and training data.
2 Related Works
Neural Surrogate Models.
Neural surrogate models are neural network-based approximations used to replace computationally expensive simulations in various scientific and engineering domains (Zhang et al., 2024; Sun and Wang, 2019). These models act as efficient emulators by learning complex input-output relationships from high-fidelity data, significantly reducing computational costs while maintaining accuracy (Raissi et al., 2020; Sun et al., 2020; Bessa et al., 2017; Thuerey et al., 2020; Raissi et al., 2019; Willard et al., 2022). Recently, generative models (e.g. pre-trained language models) have been incorporated into surrogate modeling. Some equip language models with traditional surrogate models to facilitate iterative optimization (Ma et al., 2024; Lyu et al., 2025), and some use generative models to realize the end-to-end surrogate process (Gruver et al., 2024; Hao et al., 2024; Wimmer and Rekabsaz, 2023; Che et al., 2024). While these studies primarily focus on natural sciences, time series, and multimodal gaming, the application of surrogate modeling to code execution, where both input and output exist in the modality of language, remains unexplored.
LLMs for Code.
LLMs are widely used in code-related tasks (Lu et al., 2021a; Zheng et al., 2023; Luo et al., 2023; Team, 2024a; Guo et al., 2024), which can be fundamentally categorized into code understanding and code generation. Code understanding tasks include code summarization Hu et al. (2018); Harer et al. (2019); Ahmad et al. (2020), bug detection Li et al. (2018b); Russell et al. (2018); Zhou et al. (2019); Chakraborty et al. (2020), duplication detection Zhang et al. (2019); Yu et al. (2019); Wang et al. (2020), code retrieval (Husain et al., 2020; Lu et al., 2021a), etc. Code generation tasks include code completion Li et al. (2018a); Parvez et al. (2018), code repair Chen et al. (2019); Chakraborty et al. (2020); Lutellier et al. (2020), test generation (Watson et al., 2020; Siddiq et al., 2024; schäfer2023empiricalevaluationusinglarge), etc. When evaluating LLMs on these tasks, some datasets provide broad evaluations across general tasks (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024; Muennighoff et al., 2023), some focus on specific dimensions such as multi-lingual capabilities (Athiwaratkun et al., 2022; Zheng et al., 2023; Cassano et al., 2022; Khan et al., 2023), competition problems (Hendrycks et al., 2021; Li et al., 2022), data science tasks (Chandel et al., 2022; Lai et al., 2023; Huang et al., 2022), repository-level understanding (Zhang et al., 2023a; Shrivastava et al., 2023; Liu et al., 2023; Li et al., 2024; Jimenez et al., 2023; Ding et al., 2024; Zan et al., 2024), and reasoning abilities (Athiwaratkun et al., 2022; Austin et al., 2021; Cobbe et al., 2021; Gao et al., 2023; Gu et al., 2024). However, while the potential execution result of code is important for both code understanding and generation, this aspect remains largely unexplored (Weber et al., 2024).
3 SURGE
To comprehensively evaluate the potential of LLMs as surrogate code executors, we construct a diverse benchmark, SURGE, that covers various code execution scenarios. Our dataset is designed to assess the model’s ability to approximate execution results across multiple dimensions, including multi-lingual diversity, repository-level complexity, computational intensity, error handling, and scenario-dependent variability. Below, we describe each component of SURGE, explaining the motivation behind its inclusion.
3.1 Dataset Construction
As illustrated in Figure 3, the construction of SURGE involves four distinct methodologies, each applied to specific components of our eight subsets. For ML, CL, BG components, we employ the Iterative Refactor methodology, where we refine the code through an interactive process involving LLM assistance and human verification. The RL component is constructed through Repository Sampling, where we extract and construct test cases from both public and custom code repositories. For SC, TC, DR components, we utilize Manual Implementation, carefully handcrafting code based on selected textbook materials and question collections. Finally, the FL component is developed using the Inference & Verification approach, leveraging formal mathematical provers to generate proofs and validate them through compiler verification. These methodologies are strategically applied to their respective components to ensure a comprehensive and rigorous benchmark construction.
3.2 Dataset Components
Multi-lingual Code (ML).
A fundamental nature of a general-purpose surrogate executor is its ability to handle multiple programming languages. Since we are exploring the model’s execution capabilities, we do not focus on rendering languages such as HTML, but concentrate on computational languages. Our dataset covers 7 such languages, including C, C++, C#, Java, Rust, Python, and Julia. Our dataset is adapted from McEval (Chai et al., 2024). The original dataset does not provide executable code, so we used an LLM to generate executable code by providing it with prompts, ground truth, and test cases in the original dataset. This generated code was then manually processed. We manually modified code that failed to compile, such as adding missing headers. And then we took precautions to prevent answers from being leaked in assert statements or comments.
Competition-level Code (CL).
Next, we consider competition-level code, which presents a higher level of coding difficulty. We collect these tasks from 2 public repositories 111https://github.com/azl397985856/leetcode222https://gitee.com/shanire/OJCode, which contain problems from open coding platforms (e.g. LeetCode 333https://leetcode.com, Luogu 444https://www.luogu.com.cn). The dataset includes problems in 3 languages, C++, Java, and JavaScript. Since the original repositories only provide partial solutions, we first use an LLM to generate complete, executable code that prints the expected output. This generated code is then manually verified. To investigate whether problem difficulty affects the performance of surrogate models, we further employ an LLM to automatically classify problems into different difficulty levels, followed by human verification to ensure accuracy.
Repository-level Code (RL).
In real-world scenarios, most code exists at the repository level, making repository-level code execution prediction equally important for a general-purpose surrogate model. We manually collect computational repositories that fit within the input length constraints of LLMs. These repositories include tasks such as solving the 24-point problem, Sudoku solving, and converting Python code to LaTeX. These tasks exhibit complex logic but do not rely on sophisticated models or external inputs. To assess the model’s ability to understand multi-file structures, we also manually construct two repositories containing advanced C++ syntax and multiple files.
Scientific Computing (SC).
Scientific computing has long been adopting neural surrogate models. It’s crucial to investigate whether LLMs can approximate execution results for these non-trivial tasks and hold the potential to serve as efficient surrogate models. We introduced tasks ranging from solving ordinary differential equations (ODEs) to optimization problems and signal processing. These tasks are motivated by and widely used in real-world scientific challenges, including areas where increasing research has been done on solving these computational tasks through building efficient surrogate models (Wu et al., 2023; Zhang et al., 2023b)..
Time-Consuming Algorithms (TC).
Surrogate models were originally motivated by real-world applications where program execution is high-cost and time-consuming. It’s a necessity for LLms to generalize well to strongly computation-power-dependent and time-consuming tasks. We include examples from linear algebra, sorting, searching, Monte Carlo simulations, and string matching programs, ensuring a broad representation of computationally intense tasks. These tasks cover various complexity classes, including P (e.g. sorting an array), NP (e.g. Hamiltonian Cycle), and NP-Hard (e.g. Traveling Salesman’s Problem). Additionally, we record the CPU execution time for each program under consistent environments individually to support subsequent studies.
Buggy Code (BG).
Real-world code execution often encounters errors due to syntax mistakes, logical flaws, or runtime exceptions. Code errors pose risks in sensitive scenarios, therefore, we aim for the code surrogate model to recognize the presence of bugs. This dataset is adapted from DebugBench (Tian et al., 2024), which extracted Java, Python, and C++ code from LeetCode and manually inserted errors from 4 major bug categories and 18 minor types. Since DebugBench only provides erroneous code snippets rather than complete executable programs, we first used an LLM to automatically complete the error-free code into fully runnable versions. After verifying their correctness, we replaced relevant parts with buggy code and executed them again to capture the corresponding error outputs. Some errors resulted in infinite loops causing timeouts, so we set a -second execution filter out such cases.
Code with Differential Results under Different Scenarios (DR).
Various contextual factors, such as compiler versions, optimization levels, and language standards often influence code execution. These variations can lead to different outputs for the same code snippet. It is crucial for surrogate models to recognize such discrepancies and adapt to different settings. We focus specifically on C++ and manually collect code snippets from textbooks and online sources (Bryant and O’Hallaron, 2010; Lippman et al., 2012) that exhibit different behaviors under varying compilation settings. We consider multiple compilers (g++, clang++), C++ standards (03, 11, 14, 17), and optimization levels (-O0, -O1, -O2, -O3, -Os). Each snippet is executed across these different settings, and we retain only those that produce varying outputs through different configurations while discarding cases that yield identical results across all settings.
Mathematics Formal Language (FL).
Math-Proving Formal Languages are specialized programming languages designed for mathematical proof verification through compilers (De Moura et al., 2015; Moura and Ullrich, 2021; Paulson, 1994; Barras et al., 1997). These compilers can determine whether a proof is correct and identify specific errors if present. Formal theorem proving is gaining increasing attention, as constructing a valid proof requires extensive trial and verification, which can be highly time-consuming. We aim to explore whether a surrogate model can assist in this verification process. Unlike conventional code execution, formal language verification follows a distinct paradigm, therefore we categorize this task separately. In this study, we focus on Lean 4 which is the most widely used proof assistant language. To build our dataset, we use state-of-the-art formal math prover, Goedel-Prover (Lin et al., 2025), to conduct large-scale reasoning on Lean-Workbook (Ying et al., 2024), extracting an equal proportion of correct and incorrect proofs. This balanced dataset allows us to evaluate the surrogate model’s ability to assess proof validity effectively.
| Subset | Construction Method | Source | Quantity | Example Num. | Metric | Categories |
| ML | Iterative Refactor | McEval (Chai et al., 2024) | 150 | 3 | Exact Match | Languages |
| CL | Iterative Refactor | GitHub | 150 | 3 | Exact Match | Difficulties & Languages |
| RL | Repository Sampling | GitHub & Custom | 60 | 3 | Mixed | - |
| SC | Manual Implementation | Custom | 150 | 3 | Mixed | Scenarios |
| TC | Manual Implementation | Custom | 150 | 3 | Mixed | Scenarios, CPU Time |
| BG | Iterative Refactor | DebugBench (Tian et al., 2024) | 150 | 4 | Jaccard similarity | Error Type & Languages |
| DR | Manual Implementation | Custom | 200 | 3 | Jaccard similarity | Variable Type |
| FL | Inference & Verification | Lean-Workbook (Ying et al., 2024) | 150 | 3 | Custom | - |
3.3 Evaluation Metrics
We design different evaluation metrics tailored to each subset of SURGE to ensure accurate evaluation.
In ML and CL, the outputs are simple numerical values or formatted strings and contain no error or warning message, we employ exact string matching to measure correctness.
For RL, we employ different evaluation methods for different tasks. For structured C repositories, we use exact character matching to compare outputs. For Sudoku and 24-point problems, we use edit distance to compare results. For other types of repositories, we apply the Ratcliff/Obershelp (Ratcliff and Metzener, 1988) algorithm, which measures the similarity between two sequences by finding the longest common subsequence and computing a similarity ratio.
For SC and TC, various tasks necessitate distinct evaluation methods. Specifically, (1) numerical simulations are evaluated using the average Relative Absolute Error (RAE), a widely used metric that measures the deviation between the estimated values and the ground truth values; (2) position-based tasks, such as binary search, are assessed through exact string matching; and (3) sorting tasks are evaluated by the rank correlation coefficient (Spearman, 1904), which quantifies the similarity in the ordering of elements.
For BG, where outputs contain error messages, we use the Jaccard similarity (Jaccard, 1901) between predicted and ground truth error messages. Jaccard similarity measures the overlap between two sets and is defined as , where and represent the sets of words in the predicted and ground truth error messages respectively. Jaccard similarity well captures keywords in strings, thus suitable for error message comparison.
For DR, since the same code can produce different outputs in varying settings, which sometimes include warnings or errors, we again utilize Jaccard similarity. This metric appropriately handles cases where the ground truth is not an error message.
For FL, the results consist of two parts: (1) whether the proof passes or not, and (2) if it fails, the associated error message. The evaluation proceeds as follows. If both predicted and ground truth results indicate a successful proof, the prediction is considered correct. If one is passed and the other is not, the prediction is incorrect. If both fail, we evaluate the accuracy of the error message. The error message consists of a list containing the error locations and descriptions. We compute the score of a prediction as , where is the number of errors in the ground truth, is the set of predicted error positions, represents the -th ground truth error position, represents the predicted error position corresponding to , is the indicator function which equals to when there exists and equals to when there’s not, is the ground truth error message for position , is the predicted error message for position , and is the Jaccard similarity function.
3.4 Dataset Statistics
Except for RL, which has 60 questions, and DR, which has 200 questions, all other subsets have 150 questions. Our dataset contains a total of 1,160 questions. Table 1 presents detailed statistics of SURGE, including the construction methods, problem sources, dataset quantities, number of examples in few-shot scenarios, evaluation metrics, and classification criteria for each subset.
| Model | ML | CL | RL | SC | TC | BG | DR | FL | Avg. | |||||||
| CPP | Rust | Python | Julia | Java | Others | CPP | Java | Python | ||||||||
| Zero-shot | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| GPT-4o-Mini | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-3B-Instruct | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-14B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-72B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-3B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-14B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| LLaMA-3.3-70B-Instruct | ||||||||||||||||
| Zero-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| GPT-4o-Mini | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-3B-Instruct | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-14B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-72B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-3B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-14B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| LLaMA-3.3-70B-Instruct | ||||||||||||||||
| Few-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| GPT-4o-Mini | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-3B-Instruct | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-14B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-72B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-3B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-14B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| LLaMA-3.3-70B-Instruct | ||||||||||||||||
4 Experiments
4.1 Setup
Models.
We tested SURGE on open-source and closed-source models of different sizes, including both chat models and code models. The closed-source models include GPT-4o (2024-08-06) (OpenAI, 2024b), GPT-4o-mini (2024-07-18) (OpenAI, 2024a), Claude-3.5-Sonnet (2024-10-22) (Anthropic, 2024a), and Qwen-Max (2025-01-25) (Team, 2024b). The open-source models include LLaMA-3.1-{8, 70}B-Instruct555https://huggingface.co/collections/meta-llama/llama-31-669fc079a0c406a149a5738f, LLaMA-3.3-70B-Instruct666https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, Qwen-2.5-{0.5, 1.5, 3, 7, 14, 32, 72}B-Instruct777https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e, Qwen-2.5-Coder-{0.5, 1.5, 3, 7, 14, 32}B-Instruct888https://huggingface.co/collections/Qwen/qwen25-coder-66eaa22e6f99801bf65b0c2f and DeepSeek-V3999https://huggingface.co/deepseek-ai/DeepSeek-V3 (671B).
Settings.
We tested the above models on SURGE under 3 settings: 0-shot w/o CoT, 0-shot w/ CoT, and few-shot w/ CoT. CoT here means whether we use Chain-of-Thought (Wei et al., 2022) prompting, allowing the models to think step by step, or ask the models to answer directly. We set the temperature to , i.e. employing greedy decoding.
4.2 Results
Table 2 presents the performance of 21 models across 8 sub-datasets of SURGE under 3 different settings. The best performance for each sub-dataset in each setting is highlighted in bold. In this table, we provide a detailed breakdown of the models’ performance on the ML and BG sub-datasets.
From the experimental results, several notable findings emerge:
SURGE demonstrates strong discriminative ability, especially in certain subsets. Even the strongest models perform only moderately well, highlighting the value of our benchmark. The models exhibit significant performance differences across different subsets, reflecting the comprehensiveness of our dataset. Additionally, models that perform well in other tasks (e.g. code generation), such as Claude-3.5-Sonnet, also achieve substantial overall results on SURGE. This demonstrates the reasonableness of our dataset and its effectiveness in benchmarking LLMs as general-purpose surrogate code executors.
Different prompting strategies lead to varying model performance and have different effects across subsets. We found that for the task of code execution surrogacy, both Chain-of-Thought prompting and few-shot learning can enhance model performance.
Larger model sizes tend to yield better performance on SURGE. From the results, we observe that regardless of whether it is a Qwen-Chat model or a Qwen-Coder model, performance improves as the parameter size increases.
Chat models and coder models exhibit different performance patterns and are affected differently by prompting strategies. We observed that for chat and code models of the same size (e.g. Qwen-2.5-32B-Instruct and Qwen-2.5-Coder-32B-Instruct), code models outperform chat models in the zero-shot setting SURGE. However, in the other two settings, chat models perform better. This suggests that code models have stronger zero-shot surrogate capabilities, whereas chat models excel in reasoning and imitation abilities.
On some sub-datasets of SURGE, stronger models perform worse than smaller, weaker ones. For example, in the FL dataset, we found that this occurs because stronger models tend to actively look for errors in the code, often misidentifying correct code as incorrect. In contrast, smaller models are more inclined to assume that all code is error-free. Since half of the samples in this subset are indeed correct, the smaller models end up achieving better performance.
5 Analysis
| Model | Compiler | Standard | Optimization | ||||||
| Zero-shot | Zero-shot CoT | Few-shot CoT | Zero-shot | Zero-shot CoT | Few-shot CoT | Zero-shot | Zero-shot CoT | Few-shot CoT | |
| Claude-3.5 | 12.92 | 18.19 | 25.70 | 14.21 | 16.75 | 23.59 | 16.06 | 1.23 | 12.90 |
| GPT-4o | 14.75 | 14.21 | 22.08 | 15.33 | 14.61 | 20.02 | 7.66 | 1.63 | 1.22 |
| LLaMA-3.1-8B | 3.97 | 10.12 | 12.05 | 4.29 | 10.60 | 13.17 | 1.96 | 3.91 | 8.04 |
| LLaMA-3.1-70B | 15.73 | 16.86 | 13.27 | 16.36 | 16.27 | 13.24 | 15.04 | 2.24 | 3.73 |
5.1 The Impact of Language Type
In Table 2, we compare model performance across different programming languages.
In the ML subset, models perform best in Python, followed by C++. Python’s simple syntax and dynamic typing make it easier to process, while C++ benefits from its strong presence in competitive programming and system-level tasks. Models also perform well in Julia, likely due to its clean syntax and similarity to Python. However, performance drops significantly in Rust, where strict ownership and lifetime rules introduce complexity, making code harder to predict.
In the BG subset, when predicting the output of buggy code, models excel in Python but struggle with C++. Python’s clear error messages and interpreted nature aid prediction, whereas C++’s static typing, manual memory management, and potential for undefined behavior make error handling more difficult.
5.2 The Impact of Problem Difficulty
To analyze how problem difficulty affects model performance, we examine results in the CL subset, as shown in Table 4.
Across all settings, model performance generally declines as problem difficulty increases. However, at the highest difficulty level, performance improves slightly. This anomaly arises because difficulty levels are categorized based on coding complexity, but the hardest problems often involve implementing simple functionality using complex, optimized algorithms. In such cases, models can generate correct answers through end-to-end reasoning without fully understanding the underlying code logic. In contrast, levels 1–4 require logical reasoning based on code execution, making prediction harder as complexity increases.
Furthermore, while Chain-of-Thought prompting improves performance, few-shot learning does not and may even degrade results. This is likely because buggy code varies widely, causing few-shot examples to mislead rather than aid the model.
| Difficulty | 1 | 2 | 3 | 4 | 5 |
| Zero-shot | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Zero-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Few-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
5.3 The Impact of Program’s CPU Execution Time
We explore the relationship between the execution time of the given code through running with CPUs and the accuracy of using large language models as surrogate models to acquire the output. As introduced in Section 3.1, we recorded the CPU execution times of the programs in TC. We categorize problems in TC into distinct bins based on their execution times. Next, we calculated the average accuracy for each model across all samples within the same bin, as shown in Figure 4. We observed the trend of prediction accuracy of the model falling as the actual execution time required for the corresponding program prolonged. It’s especially worth noting that for computational tasks that require execution time longer than 1 second, even state-of-the-art models struggle to obtain even one correct answer.
5.4 The Impact of Variable Type
The DR subset in SURGE examines model performance in predicting code behavior under different environmental factors. Specifically, DR includes variations in C++ compiler version, C++ standard, and compilation optimization settings. Table 3 detailedly presents model performance across different prompting strategies in DR.
In the zero-shot setting, model accuracy improves sequentially across the three factors, suggesting that compiler version differences are harder to predict than standard variations, which in turn are harder than optimization settings. However, with Chain-of-Thought reasoning, performance declines across all factors, with the sharpest drop for optimization settings. This indicates that while CoT aids reasoning for compiler versions and standards, it adds unnecessary complexity for optimizations, ultimately reducing accuracy.
5.5 Error Analysis
To further understand the model’s performance and limitations regarding serving as general surrogate models, we categorize the Errors made by Claude-3.5-Sonnet, GPT-4o and Llama-3.1-8B-Instruct and Llama-3.1-70B-Instructin the CL subset of SURGE.
We first combined machine-assisted annotation and manual verification to identify most typical error types: 1. Code Language Knowledge: evaluating foundational programming language proficiency, 2. Context Awareness: measuring understanding of long text and repository-level code, 3. Conceptual Understanding: assessing comprehension of programming concepts, 4. Calculation Process: verifying computational step accuracy, 5. Code Understanding: testing comprehension of code logic and structure, 6. Data Interpretation: evaluating data processing and analysis capabilities, and 7. Calculation Accuracy: measuring precision in scientific computations. We then use LLM to label errors in the model’s responses. The categorized error statistics are demonstrated in Figure 5.
As models prompted with CoT exhibit fewer instances of most error types compared to zero-shot, especially for the Code Understanding capability of Llama, which shows a significant improvement over zero-shot, CoT prompting results in a universal performance gain on the overall performance. The main error types are Code Understanding, Calculation Process, and Calculation Accuracy. For zero-shot, the primary error is accuracy, but for CoT, the most frequent error is Calculation Process. This suggests that CoT can better grasp the overall code logic and produce more correct results, but it may still make mistakes in the chain of thought process. In general, CoT has fewer and smaller errors. From the model perspective, Llama has a clear lead in Conceptual Understanding errors, indicating its weaker ability to understand concepts. The Claude model has the fewest errors, showing better performance under our criteria. Moreover, the error distribution of Claude and GPT is quite similar, which may suggest they share a similar way of thinking.
5.6 Training Scale Analysis
Apart from benchmarking, we systematically investigate whether training scaling can affect LLMs’ surrogate execution capabilities, we conducted comprehensive scaling experiments by training models of varying sizes on different amounts of training data. We conducted comprehensive scaling experiments with a dataset of 260K formal proofs and a fixed training batch size of 128.
As demonstrated in Figure 2, both model size and training steps are crucial factors in determining surrogate execution accuracy. As we scale from 0.5B to 7B parameters, models consistently show improved learning efficiency and higher performance ceilings throughout the training process. Larger models not only learn faster in the early stages but also continue to improve for longer before plateauing, suggesting better utilization of the training data.
These empirical observations align with established scaling laws in language modeling, indicating that surrogate execution capabilities follow similar scaling patterns as other language tasks. The consistent performance improvements with increased model size and training steps suggest that larger language models have significant potential to evolve into general-purpose surrogate code executors.
6 Conclusion
In this paper, we introduce SURGE, a holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. The curated dataset spans multiple domains, bridging theoretical tasks with real-world applications. Through extensive empirical study, we argue that while current large language models possess the power to predict code execution results to a certain extent, there remains significant room for further improvements on grounding LLMs to facilitate general surrogate model.
Acknowledgements
This work was independently conducted by the authors and is self-funded by BHL without institutional affiliation.