### v4

# SURGE: On the Potential of Large Language Models as General-Purpose Surrogate Code Executors

Abstract
Neural surrogate models are powerful and efficient tools in data mining. Meanwhile, large language models (LLMs) have demonstrated remarkable capabilities in code-related tasks, such as generation and understanding. However, an equally important yet underexplored question is whether LLMs can serve as surrogate models for code execution prediction. To systematically investigate it, we introduce SURGE, a comprehensive benchmark with problems covering key aspects: multi-language programming tasks, competition-level programming problems, repository-level code analysis, high-cost scientific computing, time-complexity-intensive algorithms, buggy code analysis, programs dependent on specific compilers or execution environments, and formal mathematical proof verification. Through extensive analysis of open-source and proprietary LLMs, we examine scaling laws, data efficiency, and predictive accuracy. Our findings reveal important insights about the feasibility of LLMs as efficient surrogates for computational processes. The benchmark and evaluation framework are available at https://github.com/Imbernoulli/SURGE.
SURGE: On the Potential of Large Language Models as General-Purpose Surrogate Code Executors
Bohan Lyu1 * † Siqiao Huang2 * Zichen Liang1 * Qi-An Sun1 Jiaming Zhang1 1Department of Computer Science and Technology, Tsinghua. 2Institute for Interdisciplinary Information Sciences (IIIS), Tsinghua. lyubh22@gmail.com, {huang-sq23, liang-zc22}@mails.tsinghua.edu.cn *Equal contribution †Corresponding author
1 Introduction
Neural surrogate models (Zhang et al., 2024; Sun and Wang, 2019) are powerful tools in data mining and machine learning, which efficiently approximate complex computational processes. Meanwhile, Large language models (LLMs) (Reid et al., 2024; Meta, 2024; Anthropic, 2024b; Hui et al., 2024; Bi et al., 2024) have demonstrated remarkable capabilities in code-related tasks (Lu et al., 2021a; Zheng et al., 2023; Luo et al., 2023; Team, 2024a; Guo et al., 2024), including code understanding (Ahmad et al., 2020; Chakraborty et al., 2020) and code generation (Li et al., 2018a; Parvez et al., 2018). However, an equally important yet underexplored question is whether LLMs can serve as general-purpose surrogate code executors, which predict the behavior of a program without actually running it. A recent study (Lyu et al., 2024) acknowledges its importance, however, it focuses on a case study rather than a systematic analysis.
The ability to predict code execution outcomes without execution has tremendous significance. In scientific computing, running simulations often requires substantial computational resources and is time-consuming, making it impractical to test every possible configuration (Lu and Ricciuto, 2019; Hesthaven and Ubbiali, 2018; Benner et al., 2015). In security-sensitive environments, executing untrusted code poses inherent risks, necessitating alternative mechanisms for assessing program behavior without exposing the system to potential vulnerabilities (Nebbione and Calzarossa, 2023; Shirazi et al., 2017; Wang et al., 2024). Additionally, some code requires highly specific execution environments, which may not always be available, making surrogate execution a valuable alternative (Queiroz et al., 2023; Gu et al., 2025). Moreover, accurately predicting a model’s potential outputs or errors is crucial for improving traditional tasks such as code understanding, code generation, and even math reasoning (Li et al., 2025). Lastly, many works use LLMs as reward models (RMs) in reinforcement learning. For code tasks, accurate execution prediction is key to a reliable RM (Ouyang et al., 2022).
Traditional approaches to surrogate code executing (King, 1976; Cadar and Sen, 2013) struggle to generalize across languages and suffer from scalability issues when applied to complex real-world codebases. Containerized environments (Merkel, 2014) mitigate dependency issues but still require full code execution. Recent efforts to train neural executors (Yan et al., 2020) focus on narrow tasks and lack the generality needed for real-world code. In contrast, LLMs’ capacity to internalize patterns from vast code corpora (Lu et al., 2021b; Chaudhary, 2023) suggests a path toward general-purpose surrogate code execution.
To understand the potential of LLMs as GEneral-purpose SURrogate code executors, we introduce SURGE. It includes components: (1) fundamental programming tasks in multiple languages, (2) competition programming problems requiring deep logical inference, (3) repository-level codebases that test long-range dependencies, (4) scientific simulations and optimizations where direct execution is high-cost, (5) time-consuming logical algorithms that have high time-complexity, (6) buggy code that examines LLMs’ ability to predict runtime errors, (7) programs whose behavior depends on specific compiler versions or execution environments and (8) math theorem proving in formal language (De Moura et al., 2015; Moura and Ullrich, 2021) which expects compilers to testify.
Through extensive evaluation of 21 open-source and proprietary LLMs on SURGE, we provide the first large-scale study of LLMs’ capabilities as computational surrogates. Additionally, we investigate the impact of various factors, including prompt engineering strategies, programming language characteristics, computational complexity, and execution time requirements, on surrogate performance. Our findings reveal both the promising potential and current limitations of LLMs as general-purpose code execution surrogates. The performance of typical models on SURGE is shown in Figure 1.
Beyond benchmarking, we conduct a scaling law study on whether LLMs’ performance improves with model size and training data. We train models with 4 different sizes on different scales of training data from the formal language subset of SURGE. Our experiments demonstrate that models’ performance consistently improves with both model size and training steps, with larger models showing stronger learning capacity and higher performance ceilings throughout the training process (Figure 2).
In short, our work makes the following key contributions:
We introduce SURGE, the first holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. It consists of subsets and problems.
We evaluate open-source and proprietary LLMs on SURGE and conduct the first large-scale analysis on them.
We present a scaling law study with models of varying sizes and scales of training data, providing empirical insights on the scaling law of LLMs on these tasks.
2 Related Works
Neural Surrogate Models.
Neural surrogate models are neural network-based approximations used to replace computationally expensive simulations in various scientific and engineering domains (Zhang et al., 2024; Sun and Wang, 2019). These models act as domain-specific emulators by learning complex input-output relationships from high-fidelity data(Raissi et al., 2020; Sun et al., 2020; Bessa et al., 2017; Thuerey et al., 2020; Raissi et al., 2019; Willard et al., 2022). Recently, generative models have been incorporated into surrogate modeling. Some equip language models with traditional surrogate models to facilitate iterative optimization (Ma et al., 2024; Lyu et al., 2025), and some use generative models to realize the end-to-end surrogate process (Gruver et al., 2024; Hao et al., 2024; Wimmer and Rekabsaz, 2023; Che et al., 2024). While these studies primarily focus on natural sciences, time series, and multimodal gaming, the application of surrogate modeling to code execution, where both input and output exist in the modality of language, remains unexplored.
LLMs for Code.
LLMs are widely used in code-related tasks (Lu et al., 2021a; Luo et al., 2023; Team, 2024a; Guo et al., 2024), which can be fundamentally categorized into code understanding and code generation. Code understanding tasks include code summarization Hu et al. (2018); Harer et al. (2019); Ahmad et al. (2020), bug detection Li et al. (2018b); Russell et al. (2018); Zhou et al. (2019); Chakraborty et al. (2020), duplication detection Zhang et al. (2019); Yu et al. (2019); Wang et al. (2020), code retrieval (Husain et al., 2020; Lu et al., 2021a), etc. Code generation tasks include code completion Li et al. (2018a); Parvez et al. (2018), code repair Chen et al. (2019); Chakraborty et al. (2020); Lutellier et al. (2020), test generation (Watson et al., 2020; Siddiq et al., 2024; Schäfer et al., 2023), etc. However, while the potential execution result of code is important for both code understanding and generation, this aspect remains largely unexplored (Weber et al., 2024).
3 SURGE
SURGE assesses the model’s ability to approximate execution results across multiple dimensions, including multi-lingual diversity, repository-level complexity, computational intensity, error handling, and scenario-dependent variability. Below, we describe each component of SURGE, including motivation and construction methods.
| Subset | Construction Method | Source | Quantity | Metric | Categories |
| ML | Iterative Refactor | McEval (Chai et al., 2024) | 150 | Exact Match | Languages |
| CL | Iterative Refactor | GitHub | 150 | Exact Match | Difficulties & Languages |
| RL | Repository Sampling | GitHub & Custom | 60 | Mixed | - |
| SC | Manual Implementation | Custom | 150 | Mixed | Scenarios |
| TC | Manual Implementation | Custom | 150 | Mixed | Scenarios, CPU Time |
| BG | Iterative Refactor | DebugBench (Tian et al., 2024) | 150 | Jaccard similarity | Error Type & Languages |
| DR | Manual Implementation | Custom | 200 | Jaccard similarity | Variable Type |
| FL | Inference & Verification | Lean-Workbook (Ying et al., 2024) | 150 | Custom | - |
3.1 Dataset Construction
As illustrated in Figure 3, the construction of SURGE involves four distinct methodologies, each applied to specific components of our eight subsets. For ML, CL, BG components, we employ the Iterative Refactor methodology, where we refine the code through an interactive process involving LLM assistance and human verification. The RL component is constructed through Repository Sampling, where we extract and construct test cases from both public and custom code repositories. For SC, TC, DR components, we utilize Manual Implementation, carefully handcrafting code based on selected textbook materials and question collections. Finally, the FL component is developed using the Inference & Verification approach, leveraging formal mathematical provers to generate proofs and validate them through compiler verification.
To mitigate the risk of data contamination and answer leakage, we implemented a robust two-fold sanitization strategy across the dataset: (1) We applied automated filtering scripts to systematically remove comments from all code snippets before presenting them to the models for evaluation. (2)We manually inspected all generated or collected code to identify and remove any comments, ‘assert‘ statements, or other annotations that could inadvertently provide hints or reveal the expected output.
3.2 Dataset Components
Multi-lingual Code (ML).
A fundamental nature of a general-purpose surrogate executor is its ability to handle multiple programming languages, especially computational languages. Our dataset covers 7 such languages, including C, C++, C#, Java, Rust, Python, and Julia. Our dataset is adapted from McEval (Chai et al., 2024). The original dataset does not provide executable code, so we used an LLM to generate executable code by providing it with prompts, ground truth, and test cases in the original dataset. This generated code was then manually processed; specifically, we (1) modified code that failed to compile (e.g., by adding missing headers) and (2) carefully reviewed the code to prevent answer leakage through assert statements or comments.
Competition-level Code (CL).
Next, we consider competition-level code, which presents a higher level of coding difficulty. We collect these tasks from 2 public repositories 111https://github.com/azl397985856/leetcode222https://gitee.com/shanire/OJCode, which contain problems from open coding platforms (e.g. LeetCode, Luogu). The dataset includes problems in 3 languages, C++, Java, and JavaScript. Since the original repositories only provide partial solutions, we first use an LLM to generate complete, executable code that prints the expected output. This generated code is then manually verified. During this process, if issues such as package mismatches or syntax errors (e.g., mixing language versions) arose, we manually revised the code to ensure successful compilation and execution. To investigate whether problem difficulty affects the performance of surrogate models, we classify problems into different difficulty levels.
Repository-level Code (RL).
In real-world scenarios, most code exists at the repository level, making repository-level code execution prediction equally important for a general-purpose surrogate model. We manually collect computational repositories that fit within the input length constraints of LLMs. These repositories include tasks such as solving the 24-point problem, Sudoku solving, and converting Python code to LaTeX. These tasks exhibit complex logic but do not rely on sophisticated models or external inputs. To assess the model’s ability to understand multi-file structures, we also manually construct two repositories containing advanced C++ syntax and multiple files.
Scientific Computing (SC).
Scientific computing has long been adopting neural surrogate models. We introduced tasks ranging from solving ordinary differential equations (ODEs) to optimization problems and signal processing. These tasks are motivated by and widely used in real-world scientific challenges, including areas where increasing research has been done on solving these computational tasks through building efficient surrogate models (Wu et al., 2023; Zhang et al., 2023). A comprehensive overview of the setup for each task, along with the corresponding algorithms, can be found in Appendix C.4.1.
Time-Consuming Algorithms (TC).
Surrogate models were originally motivated by real-world applications where program execution is high-cost and time-consuming. It’s a necessity for LLMs to generalize well to strongly computation-power-dependent and time-consuming tasks. We include examples from linear algebra, sorting, searching, Monte Carlo simulations, and string matching programs, ensuring a broad representation of computationally intense tasks. These tasks cover various complexity classes, including P (e.g., sorting an array), NP (e.g., Hamiltonian Cycle), and NP-Hard (e.g., Traveling Salesman’s Problem). Additionally, we record the CPU execution time for each program under consistent environments individually to support subsequent studies. A detailed description of this subset is provided in Appendix C.5.1.
Buggy Code (BG).
Real-world code execution often encounters errors, which pose risks in sensitive scenarios. Therefore, code surrogate models should recognize the presence of bugs. This dataset is adapted from DebugBench (Tian et al., 2024), which extracted Java, Python, and C++ code from LeetCode and manually inserted errors from 4 major bug categories and 18 minor types. Since DebugBench only provides code snippets rather than complete executable programs, we first used an LLM to automatically complete the error-free code into fully runnable versions. After verifying their correctness, we replaced relevant parts with buggy code and executed them again to capture the corresponding error outputs. Some errors resulted in infinite loops, causing timeouts, so we set a -second execution filter to avoid such cases.
Code with Differential Results under Different Scenarios (DR).
Various contextual factors, such as compiler versions, optimization levels, and language standards, often influence code execution. These variations can lead to different outputs for the same code snippet. We focus specifically on C++ and manually collect code snippets from textbooks and online sources (Bryant and O’Hallaron, 2010; Lippman et al., 2012) that exhibit different behaviors under varying compilation settings. We consider multiple compilers (g++, clang++), C++ standards (03, 11, 14, 17), and optimization levels (-O0, -O1, -O2, -O3, -Os). Each snippet is executed across these different settings, and we retain only those that produce varying outputs through different configurations while discarding cases that yield identical results across all settings.
Mathematics Formal Language (FL).
Math-Proving Formal Languages are specialized programming languages designed for mathematical proof verification through compilers (De Moura et al., 2015; Moura and Ullrich, 2021; Paulson, 1994; Barras et al., 1997). These compilers can determine whether a proof is correct and identify specific errors. The verification is very time-consuming. In this study, we focus on Lean4 which is the most widely used proof language. To build our dataset, we use Goedel-Prover (Lin et al., 2025) to conduct large-scale reasoning on Lean-Workbook (Ying et al., 2024) and extract an equal proportion of correct and incorrect proofs. This balanced dataset allows us to evaluate the surrogate model’s ability to assess proof validity effectively.
3.3 Evaluation Metrics
We design different evaluation metrics tailored to each subset of SURGE to ensure accurate evaluation.
In ML and CL, the outputs are simple numerical values or formatted strings, we employ exact string matching to measure correctness.
For RL, we employ different evaluation methods for different tasks. For structured C repositories, we use exact character matching to compare outputs. For Sudoku and 24-point problems, we use edit distance to compare results. For other types of repositories, we apply the Ratcliff/Obershelp (Ratcliff and Metzener, 1988) algorithm.
For SC and TC, various tasks necessitate distinct evaluation methods. Specifically, (1) numerical simulations are evaluated using the average Relative Absolute Error (RAE); (2) position-based tasks, such as binary search, are assessed through exact string matching; and (3) sorting tasks are evaluated by the rank correlation coefficient (Spearman, 1904). Details regarding the evaluation metrics can be found in Appendix C.4.2 and Appendix C.5.2.
For BG, we use the Jaccard similarity (Jaccard, 1901) between predicted and ground truth error messages. For DR, since the same code can produce different outputs in varying settings, which sometimes include warnings or errors, we again utilize Jaccard similarity.
For FL, the results consist of two parts: (1) whether the proof passes or not, and (2) if it fails, we evaluate the accuracy of the error message. The error message consists of a list containing the error locations and descriptions. We compute the score of a prediction as , where is the number of errors in the ground truth, is the set of predicted error positions, represents the -th ground truth error position, represents the predicted error position corresponding to , is the indicator function which equals to only when there exists , is the ground truth error message for position , is the predicted error message for position , and is the Jaccard similarity function.
| Model | ML | CL | RL | SC | TC | BG | DR | FL | Avg. | |||||||
| CPP | Rust | Python | Julia | Java | Others | CPP | Java | Python | ||||||||
| Zero-shot | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| Zero-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| Few-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
3.4 Dataset Statistics
Table 1 presents detailed statistics of SURGE, including the construction methods, problem sources, dataset quantities, number of examples in few-shot scenarios, evaluation metrics, and classification criteria for each subset.
4 Experiments
4.1 Setup
Models.
We tested SURGE on open-source and closed-source models of different sizes, including both chat models and code models. The closed-source models include GPT-4o (2024-08-06) (OpenAI, 2024b), GPT-4o-mini (2024-07-18) (OpenAI, 2024a), Claude-3.5-Sonnet (2024-10-22) (Anthropic, 2024a), and Qwen-Max (2025-01-25) (Team, 2024b). The open-source models include LLaMA-3.1-{8, 70}B-Instruct, LLaMA-3.3-70B-Instruct, Qwen-2.5-{0.5, 1.5, 3, 7, 14, 32, 72}B-Instruct, Qwen-2.5-Coder-{0.5, 1.5, 3, 7, 14, 32}B-Instruct and DeepSeek-V3 (671B).
Settings.
We tested the above models on SURGE under 3 settings: 0-shot w/o CoT, 0-shot w/ CoT, and few-shot w/ CoT. CoT here means whether we use Chain-of-Thought (Wei et al., 2022) prompting, allowing the models to think step by step, or ask the models to answer directly. We set the temperature to , i.e., employing greedy decoding.
| Model | Compiler | Standard | Optimization | ||||||
| Zero-shot | Zero-shot CoT | Few-shot CoT | Zero-shot | Zero-shot CoT | Few-shot CoT | Zero-shot | Zero-shot CoT | Few-shot CoT | |
| Claude-3.5 | 12.92 | 18.19 | 25.70 | 14.21 | 16.75 | 23.59 | 16.06 | 1.23 | 12.90 |
| GPT-4o | 14.75 | 14.21 | 22.08 | 15.33 | 14.61 | 20.02 | 7.66 | 1.63 | 1.22 |
| LLaMA-3.1-8B | 3.97 | 10.12 | 12.05 | 4.29 | 10.60 | 13.17 | 1.96 | 3.91 | 8.04 |
| LLaMA-3.1-70B | 15.73 | 16.86 | 13.27 | 16.36 | 16.27 | 13.24 | 15.04 | 2.24 | 3.73 |
4.2 Results
Table 2 presents the performance of 10 selected models across 8 sub-datasets of SURGE under 3 different settings. In this table, we provide a detailed breakdown of the models’ performance on the ML and BG sub-datasets. We present the complete results of all 21 models in Table 5 in Appendix A. From the results, several notable findings emerge:
SURGE demonstrates strong discriminative ability. Even the strongest models perform only moderately well, highlighting the value of our benchmark. The models exhibit significant performance differences across different subsets, reflecting the comprehensiveness of our dataset. Additionally, models that perform well in other tasks, such as Claude-3.5-Sonnet, also achieve substantial overall results on SURGE. This demonstrates the reasonableness of our dataset and its effectiveness in benchmarking LLMs as general-purpose surrogate code executors.
Different prompting strategies lead to varying model performance and have different effects across subsets. We found that for the task of code execution surrogacy, both Chain-of-Thought prompting and few-shot learning can enhance model performance.
Larger model sizes tend to yield better performance on SURGE. From the results, we observe that regardless of whether it is a Qwen-Chat model or a Qwen-Coder model, performance improves as the parameter size increases.
Chat models and coder models exhibit different performance patterns and are affected differently by prompting strategies. We observed that for chat and code models of the same size, code models outperform chat models in the zero-shot setting SURGE. However, in the other two settings, chat models perform better. This suggests that code models have stronger zero-shot surrogate capabilities, whereas chat models excel in reasoning and imitation abilities.
On some sub-datasets of SURGE, stronger models perform worse than smaller, weaker ones. For example, in the FL dataset, we found that this occurs because stronger models tend to actively look for errors in the code, often misidentifying correct code as incorrect. In contrast, smaller models are more inclined to assume that all code is error-free. Since half of the samples in this subset are indeed correct, the smaller models end up achieving better performance.
4.3 Anomaly Analysis
We conducted case studies on anomalous results to better understand model behavior.
Larger Models Underperforming. In some instances, larger models performed worse than their smaller counterparts. For example, on ML tasks, Qwen-2.5-14B-Instruct was outperformed by Qwen-2.5-0.5B-Instruct. Our case study revealed that while the 0.5B model’s reasoning was less detailed, it correctly grasped the code’s core logic. In contrast, the 14B model attempted more elaborate reasoning but often made critical misjudgments (e.g., in conditionals like ), leading to incorrect results. This suggests that on certain complex reasoning tasks, larger models may be prone to "overthinking" or following incorrect logical paths.
Zero-Shot Failures. We investigated cases of complete failure in the zero-shot setting, such as LLaMA-3.1-8B-Instruct scoring 0 on ML tasks for C++ and Rust, and on all FL tasks. For ML, the failures were attributed to the inherent difficulty of the test cases and specific language features (e.g., Rust’s strict type system). For FL tasks, the model incorrectly identified errors in all problems. For half of the correct problems, it wrongly predicted errors. For the other half that contained errors, it failed to predict the specific error messages accurately, leading to incorrect predictions in all cases.
5 Analysis
5.1 The Impact of Language Type
In Table 2, we compare model performance across different programming languages.
In the ML subset, LLMs perform best in Python, followed by C++. Python’s simple syntax makes it easier to process, while C++ benefits from its strong presence in programming and system-level tasks. Models also perform well in Julia, likely due to its clean syntax and similarity to Python. However, performance drops significantly in Rust, where strict ownership and lifetime rules introduce complexity, making it hard to predict.
In the BG subset, when predicting the output of buggy code, models excel in Python but struggle with C++. Python’s clear error messages aid prediction, whereas C++’s static typing, manual memory management, and potential for undefined behavior make error handling more difficult.
| Difficulty | 1 | 2 | 3 | 4 | 5 |
| Zero-shot | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Zero-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Few-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
5.2 The Impact of Problem Difficulty
To analyze how problem difficulty affects model performance, we examine results in the CL subset, as shown in Table 4.
Across all settings, model performance generally declines as problem difficulty increases. However, at the highest difficulty level, performance improves slightly. This anomaly arises because difficulty levels are categorized based on coding complexity, but the hardest problems often involve implementing simple functionality using complex, optimized algorithms. In such cases, models can generate correct answers through end-to-end reasoning without fully understanding the underlying code logic. In contrast, levels 1–4 require logical reasoning based on code execution, making prediction harder as complexity increases.
Furthermore, while Chain-of-Thought prompting improves performance, few-shot learning does not and may even degrade results. This is likely because buggy code varies widely, causing few-shot examples to mislead rather than aid the model.
5.3 Prediction Accuracy & CPU Time
We explore the relationship between the execution time of the given code through running on CPUs and the accuracy of using LLMs as surrogate models to acquire the output. We categorize problems in TC into distinct bins based on their execution times and calculate the average accuracy for each model across all samples within the same bin, as shown in Figure 4. We observed the trend of prediction accuracy of the model falling as the actual execution time required for the corresponding program prolonged. It’s especially worth noting that for computational tasks that require execution time longer than 1 second, even state-of-the-art models struggle to obtain even one correct answer.
5.4 The Impact of Variable Type
The DR subset in SURGE examines model performance in predicting code behavior under different environmental factors. Specifically, DR includes variations in C++ compiler version, C++ standard, and compilation optimization settings. Table 3 presents model performance across different prompting strategies in DR.
In the zero-shot setting, model accuracy improves sequentially across the three factors, suggesting that compiler version differences are harder to predict than standard variations, which in turn are harder than optimization settings. However, with Chain-of-Thought reasoning, performance declines across all factors, with the sharpest drop for optimization settings. This indicates that while CoT aids reasoning for compiler versions and standards, it adds unnecessary complexity for optimizations, ultimately reducing accuracy.
5.5 In-depth Look at Model Behavior
We looked into LLMs’ surrogation process and summarized three key observations:
Deep Semantic and Logical Reasoning. LLMs demonstrate a strong ability to interpret complex, language-specific semantics. For instance, when analyzing C++ code with bit fields, models correctly reasoned about the effects of signed vs. unsigned types, bit-width limitations, and how value overflows are handled under two’s complement representation. This indicates an understanding of low-level computational logic.
Stateful Algorithmic Tracing. For algorithmic tasks, LLMs effectively simulate the execution process by tracing the program’s state. The models’ chain-of-thought process shows them iterating through loops, updating variables step-by-step, and applying conditional logic as an interpreter would.
Identification of Computational Boundaries. In a brute-force Traveling Salesman Problem (TSP) solver, models could correctly understand the overall goal and algorithm, but often failed to predict the precise final floating-point value. This suggests their "execution" is a form of sophisticated logical inference, which falters when faced with complex combinatorial calculations that exceed their internal reasoning capacity.
5.6 Error Analysis
To further understand the model’s performance and limitations regarding serving as general surrogate models, we categorize the Errors made by Claude-3.5-Sonnet, GPT-4o, and Llama-3.1-8B-Instruct and Llama-3.1-70B-Instructin the CL subset of SURGE.
We first combined LLM-assisted annotation and manual verification to identify most typical error types: 1. Code Language Knowledge: evaluating foundational programming language proficiency, 2. Context Awareness: measuring understanding of long text and repository-level code, 3. Conceptual Understanding: assessing comprehension of programming concepts, 4. Calculation Process: verifying computational step accuracy, 5. Code Understanding: testing comprehension of code logic and structure, 6. Data Interpretation: evaluating data processing and analysis capabilities, and 7. Calculation Accuracy: Measuring precision in scientific computations. We then use LLM to label errors in the model’s responses. The categorized error statistics are demonstrated in Figure 5.
As models prompted with CoT exhibit fewer instances of most error types compared to zero-shot, especially for the Code Understanding capability of Llama. The main error types are Code Understanding, Calculation Process, and Calculation Accuracy. For zero-shot, the primary error is accuracy, but for CoT, the most frequent error is Calculation Process. This suggests that CoT can better grasp the overall code logic and produce more correct results, but it may still make mistakes in the chain of thought process. In general, CoT has fewer and smaller errors. From the model perspective, Llama has a clear lead in Conceptual Understanding errors, indicating its weaker ability to understand concepts.
5.7 Training Scale Analysis
We also investigate whether training scaling can affect LLMs’ surrogate execution capabilities on the FL task. We trained models of varying sizes on different amounts of training data sampled from the same distribution, and tested them to predict the error feedbacks for incorrect proofs.
As demonstrated in Figure 2, both model size and training steps are crucial factors in determining surrogate execution accuracy. As we scale from 0.5B to 7B parameters, models consistently show improved learning efficiency and higher performance ceilings throughout the training process. Larger models learn faster in the early stages and continue to improve for longer before plateauing, suggesting better utilization of the training data.
These empirical observations align with established scaling laws in language modeling, indicating that surrogate execution capabilities follow similar scaling patterns as other language tasks.
6 Conclusion
We introduce SURGE, a holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. Through extensive empirical study, we argue that there remains significant room for further improvements on grounding LLMs to facilitate general surrogate models. Our findings not only chart the current landscape but also illuminate a clear path for future research.
Acknowledgements
This work was independently conducted by the authors and is self-funded by BL without institutional affiliation.
Limitations
Despite its comprehensive evaluation, our study has several limitations. LLMs remain approximators rather than exact code executors, often struggling with edge cases, intricate runtime behaviors, and execution-dependent state changes. While SURGE covers diverse execution scenarios, it does not encompass all specialized environments, such as hardware-dependent simulations or real-time systems. Additionally, LLMs may generate plausible but incorrect outputs, particularly in complex logical dependencies or undefined behaviors, making error detection challenging. Our scaling study is constrained by computational resources, limiting the assessment of extremely large models or extensive training data distributions. Furthermore, security risks remain, as LLMs may fail to recognize vulnerabilities, potentially misjudging harmful code. Finally, our benchmark operates in a controlled setting, whereas real-world software development involves dynamic interactions and iterative debugging, which are not fully captured in our study. Future work should focus on improving LLMs’ reasoning abilities, enhancing robustness in execution prediction, and integrating them with traditional program analysis techniques for practical deployment.
Appendix
Appendix A Complete Main Results
| Model | ML | CL | RL | SC | TC | BG | DR | FL | Avg. | |||||||
| CPP | Rust | Python | Julia | Java | Others | CPP | Java | Python | ||||||||
| Zero-shot | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| GPT-4o-Mini | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-3B-Instruct | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-14B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-72B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-3B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-14B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| LLaMA-3.3-70B-Instruct | ||||||||||||||||
| Zero-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| GPT-4o-Mini | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-3B-Instruct | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-14B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-72B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-3B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-14B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| LLaMA-3.3-70B-Instruct | ||||||||||||||||
| Few-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| GPT-4o-Mini | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-3B-Instruct | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-14B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-72B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-0.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-1.5B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-3B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-14B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| LLaMA-3.3-70B-Instruct | ||||||||||||||||
Appendix B Prompts
B.1 Prompts for Dataset Refactoring
ML:
CL:
BG:
Appendix C Details of SURGE
C.1 ML
| Java | C# | Rust | Julia | Python | C++ | C |
| 25 | 20 | 20 | 26 | 18 | 21 | 20 |
In ML, the usage distribution of various programming languages is shown in Table 6. We selected a variety of languages, including Java, C#, Rust, Julia, Python, C++, and C, to evaluate the model’s ability to handle multilingual code. This diverse selection helps to comprehensively assess the model’s performance across different languages.
C.1.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.1.2 Demo Questions
C.2 CL
| Python | C++ | JavaScript |
| 50 | 51 | 49 |
| Difficulty | JavaScript | CPP | Python |
| 1 | 10 | 11 | 11 |
| 2 | 6 | 4 | 6 |
| 3 | 12 | 14 | 12 |
| 4 | 8 | 8 | 9 |
| 5 | 13 | 14 | 12 |
In CL, we selected competition problems of varying difficulty, each with solutions in Python, C++, and JavaScript. You can see the distribution of language in Table 7, and the distribution of problem difficulty in Table 8. This selection allows us to test the model’s cross-language capabilities and its ability to handle problems of different difficulty levels.
C.2.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.2.2 Demo Questions
C.3 RL
| Python | C++ |
| 24 | 36 |
In RL, the distribution of programming language usage is shown in Table 9. We utilized five GitHub repositories for this study, consisting of two Python projects and three C++ projects. Each repository contains a set of ten or more test cases, providing a diverse set of data for evaluation across different programming languages.
C.3.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.3.2 Demo Questions
C.4 SC
C.4.1 Tasks Descriptions
The scientific computing component of SURGE consists of 4 carefully curated areas, aiming to evaluate model performance on computational tasks that exhibit a time-consuming nature as well as applicational values in scientific computing areas. In this section, we provide a detailed description of each component.
Numerical Optimization.
In this task, the model is given a program that solves an optimization problem through gradient descent. The query may be the optimized value (min) or the optimal point (argmin). We carefully select four functions, which consist of: a simple quadratic function, Rosenbrock Function, Himmelblau’s Function, and a polynomial function with linear constraints. For each function, we will select multiple different hyperparameter configurations to assess the model’s performance. These four functions provide a systematic evaluation of the model’s potential to serve as a surrogate model in this field. As the quadratic function is solvable without need the to run the gradient descent, the model may solve it through world knowledge. The Rosenbrock function is known for its narrow, curved valley containing the global minimum, making it difficult for optimization algorithms to converge. Therefore the output is highly dependent on hyperparameters (initial point, learning rate, maximum steps), thus the model must execute code in its reasoning process to acquire the answer. Himmelblau’s function has multiple local minima, also posing sensitivity to hyperparameters.
PDE Solving.
We consider three types of Partial Differential Equations: the 1D Heat Equation, the 2D Wave Equation, and the 2D Laplace Equation. For the 1D Heat Equation, we focus on solving the following equation:
For the 2D Laplace Equation, we aim to solve the equation:
Lastly, for the 2D Wave Equation, we work on solving the following equation:
We solve 1D Heat Equation and 2D Wave Equation using the Explicit Finite Difference Method. For the 2D Laplace Equation, we solve it using the Gauss-Seidel Method. The model is then queried on the values of and .
Fourier Transform (FFT)
We implement FFT using the Cooley-Tukey Algorithm and query the model to give the magnitude of the top 10 values.
ODE Solving
For solving ordinary differential equations, we constructed three different equations and implemented the Euler Method and the Runge-Kutta Method so solve these equations.
C.4.2 Evaluation Metrics
Relative Absolute Error (RAE).
Given a scalar ground truth value and a model prediction , the Relative Absolute Error (RAE) is defined as:
For cases involving multiple entries, such as tensors or vectors, the following alignment procedure is applied: (1) if the prediction contains fewer elements than the ground truth, the prediction is padded with zeros until it matches the length of the ground truth; (2) if the prediction has more elements than the ground truth, it is truncated to match the ground truth length. The average RAE is then computed by averaging the RAE for each corresponding element.
Exact Matching.
For tasks involving position-based predictions, such as binary search, we adapt exact matching, as the accuracy of the algorithm is determined by comparing the exactness of the estimated result to the true result. This evaluation method checks if the estimated solution matches the ground truth exactly, typically using string or sequence matching. For such tasks, an exact match is considered a success, and any discrepancy between the ground truth and the estimate results in failure. Formally, given a string and the model’s prediction , the Exact Matching is given by:
where is the indicator function.
C.4.3 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.4.4 Demo Questions
C.5 TC
C.5.1 Tasks Descriptions of Time Consuming (TC)
The time consuming component of SURGE is comprised of 4 tasks in for computationally expensive areas, covering a spectrum of Linear Algebra, Sorting, Searching, Monte Carlo Simulations and String Matching Programs. Some of these tasks take hours to complete, showing their potential to benchmark LLM’s ability to reason through lengthy computations.
Linear Algebra.
In this task, we are focused on acquiring key properties in linear algebra given square matrices of varying sizes. In particular, we query the model on solving LU decomposition, QR decomposition, the largest eigenvalue and eigenvector using the power method, and the inversion matrix.
Sorting And Searching.
We include four classical algorithmic problems in this area, namely Hamiltonian Cycle, Traveling Salesman Problem (TSP), Sorting an array of real numbers and Searching. For Hamiltonian Cycle, we adopt the backtracking algorithm. Specifically, we randomly generate graphs with vertices from 4 to 100 and ask the model to find whether a Hamiltonian cycle exists. For TSP, we implement a naive brute-force algorithm and ask the model to find the length of the optimal path. For Sorting, we adopt the bubble sort, quick sort, and merge sort algorithms. For each algorithm, we consider different list sizes from 5 to 100 and generate 10 test cases for each list size. The evaluation metric is the rank correlation (also Spearman’s ). Lastly, for searching, we adopt binary search and query the model on randomly generated lists of varying sizes.
Monte Carlo Estimation.
We adopt Monte Carlo simulation to estimate the values of specific real numbers (e.g. ), as well as a future stock price prediction that follows the Brownian motion. We alter the number of samples used in Monte Carlo estimation, resulting in varying program outcomes.
String Matching Program.
We adopt the naive string matching, KMP, and Rabin-Karp algorithms. For each algorithm, we randomly generate text and pattern with varying lengths, and query the model on the existence and position of the matching.
C.5.2 Evaluation Metrics
Rank Correlation.
Rank Correlation (Spearman, 1904), also referred to as Spearman’s , is used to assess sorting tasks by measuring the correlation between the estimated ordinal ranking and the ground truth, which can be written as:
where and denote the true and estimated rankings, respectively, and Cov and represent the covariance and standard deviation of the respective sequences.
C.5.3 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.5.4 Demo Questions
C.6 BG
| Java | Python | C++ |
| 51 | 45 | 54 |
| Error Type | Java | Python3 | CPP |
| == and = confusion | 5 | 6 | 5 |
| undefined keywords | 6 | 3 | 5 |
| parentheses mismatch | 5 | 5 | 6 |
| indexing error | 10 | 9 | 11 |
| undefined objects | 11 | 9 | 8 |
| unclosed string | 7 | 5 | 7 |
| conditional statement error | 10 | 8 | 9 |
| undefined methods | 8 | 3 | 6 |
| colon missing | 5 | 7 | 8 |
| wrong comment mark | 9 | 1 | 9 |
| variable value error | 2 | 2 | 4 |
| operation error | 2 | 2 | 3 |
| other error | 4 | 2 | 1 |
| statement separation | 4 | 0 | 7 |
| indentation error | 0 | 4 | 0 |
| Double Bugs | 10 | 8 | 10 |
| Triple Bugs | 12 | 10 | 11 |
| Quadruple Bugs | 8 | 5 | 9 |
In BG, the distribution of language usage across categories is shown in Table 10, indicating a balanced usage of Java, Python, and C++. Table 11 presents a detailed breakdown of bug types and their frequency across different languages. This distribution allows us to assess the model’s ability to handle a variety of bugs across multiple programming languages.
C.6.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.6.2 Demo Questions
C.7 DR
C.7.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.7.2 Demo Questions
C.8 FL
Hard for Latex to Compile, check the source code.
Appendix D Training Details
For training, we employ Llama-Factory (Zheng et al., 2024) as the LLM training platform. Table 12 shows our training hyperparameters.
| Parameter | Value |
| Train batch size | 128 |
| Learning rate | 1.0e-5 |
| Number of epochs | 2.0 |
| LR scheduler | cosine |
| Warmup ratio | 0.1 |
| Precision | bf16 |